# Value Residual Learning

### v4

@@ 4.6 Post-Analysis of ResFormer
Sequence Length Model Valid Loss
@@
Sequence Length Model Valid Loss
@@
Architecture Learning Rate Warnup Steps Valid Loss
Llama (82M) 1e-4 120 -0.033
@@ A.1 Attention pattern analysis
A.2 Downstream Evaluations
We compare the ResFormer and vanilla Transformer on several classical reasoning tasks following (Zhang et al., 2024) in a zero-shot way. The tasks include Hellaswag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2019), ARC-Easy and ARC-Challenge (Clark et al., 2018) and PIQA (Bisk et al., 2020). The results in Table 11 show that ResFormer achieved an average accuracy improvement of nearly 3% compared to the vanilla Transformer.
Model Max Length HellaSwag Obqa WinoGrande ARC-c ARC-e PIQA Avg
Transformer 2,048 0.263 0.142 0.492 0.199 0.331 0.572 0.333
ResFormer 2,048 0.273 0.148 0.512 0.182 0.414 0.604 0.355
Transformer 64,000 0.267 0.142 0.485 0.179 0.322 0.570 0.328
ResFormer 64,000 0.274 0.136 0.513 0.184 0.407 0.588 0.350
Table 11: Zero-shot accuracy on commonsense reasoning tasks.
A.3 Pre-train Dataset
@@ A.4 Training Details
Section 4.1 introduces the main experimental hyperparameters used in the paper. This section further details the training parameters for various model sizes and training sequence lengths. Table 14 demonstrates the differences among models of various sizes. The configurations for the number of layers, attention heads, hidden dimensions, and FFN dimensions are based on Biderman et al. (2023). Additionally, the $\bm{\lambda}$ in Eqn. 3 is set to be 0.4 for NeuTRENO. Moreover, as reported in Table 13, the batch size that a single GPU can accommodate varies depending on the length of the training sequences. Note that the total number of tokens in each batch is consistently 2 million.

### v5

@@ 4.5 Post-Analysis of ResFormer
Sequence Model Valid
Length Loss
@@
Sequence Model Valid
Length Loss
@@
Model Type Learning Rate Warnup Steps Valid Loss
Llama 1e-4 120 -0.033
@@ 5 Conclusion
Acknowledgments
This work was supported by the Scientific Research Project of Westlake University (Grant No. WU2024B003).
@@ A.1 Attention pattern analysis
A.2 Pre-train Dataset
@@ A.3 Training Details
Section 4.1 introduces the main experimental hyperparameters used in the paper. This section further details the training parameters for various model sizes and training sequence lengths. Table 14 demonstrates the differences among models of various sizes. The configurations for the number of layers, attention heads, hidden dimensions, and FFN dimensions are based on Biderman et al. (2023). Moreover, as reported in Table 13, the batch size that a single GPU can accommodate varies depending on the length of the training sequences. Note that the total number of tokens in each batch is consistently 2 million.